# Your AI, Not Your View: The Bias of LLMs in Investment Analysis

### v1

@@ 4.3. Decision Uncertainty
$H(\text{Decision})=-\sum_{x\in\{\text{buy, sell}\}}P(x)\log_{2}P(x)$
A higher entropy value indicates greater uncertainty, while a lower entropy corresponds to higher confidence in the decision-making process. This analysis compares the uncertainty of two models: DeepSeek-V3 as a representative of models with overall high preference, and GPT-4.1 as a representative of models with low preference.
Figure 6. Entropy comparison between a high-preference model (DeepSeek-V3) and a low-preference model (GPT-4.1). The pattern inverts when the prompt shifts from balanced to imbalanced, with DeepSeek-V3’s confidence turning into higher entropy while GPT-4.1’s uncertainty decreases.
Under the Balanced Prompt condition, where evidence was presented in equilibrium, the two models exhibited distinct entropy patterns. GPT-4.1, characterized by its weak inherent bias, recorded high entropy, signifying a state of high uncertainty and an inability to commit to a decision. Conversely, DeepSeek-V3, with its strong initial preference, showed very low entropy. This suggests its internal bias easily broke the tie presented by the external evidence, allowing it to make a confident decision.
Interestingly, under the Imbalanced Prompt condition, where more counter-evidence was presented, the entropy pattern inverted. The entropy of DeepSeek-V3 rose sharply, suggesting it was experiencing cognitive dissonance from the conflict between its strong internal preference and the clear external counter-evidence. In contrast, the entropy of GPT-4.1 decreased. With less pre-existing bias, it could confidently align with the majority evidence, which resolved its uncertainty from the previous condition.
Ultimately, stronger inherent preferences appear to amplify hesitation and uncertainty when challenged by conflicting external evidence. The entropy analysis thus underscores how intrinsic preferences significantly influence not only the direction of decisions but also the confidence levels and internal cognitive conflict experienced by models during decision-making.
@@ 5. Limitations
This study’s limitations are as follows. First, all evidence was generated by a specific LLM, and its intensity was simplified to a single numerical value. This approach cannot fully capture the potential biases of the generator model or the richness of real-world information. Second, the current experimental design, based on differences in the volume or intensity of evidence, has limitations for evaluating reasoning models. This is because these models, rather than experiencing the conflict between contradictory information intended by the experiment, can objectively compare the given numerical values to calculate an optimal answer, making their decision a result of computational ability rather than bias. Third, our analysis is static and does not capture the temporal dynamics of model biases; it provides a snapshot at a single point in time without investigating how these preferences might change over different periods.
@@ 6. Conclusion
This study systematically investigated the intrinsic preferences of LLMs in financial contexts and analyzed how these preferences harden into entrenched biases under informational conflict. We sought to answer two key research questions regarding the intrinsic preferences LLMs hold for financial factors and how these lead to biases. The results show that LLMs are not neutral decision-makers, with distinct preferences for certain financial factors depending on the model. While sector preferences varied significantly across models, showing no overall trend, a common bias towards large-size stocks and a consistent preference for a contrarian investment view over momentum were observed.
@@
Bias verification experiments clearly revealed that these latent preferences directly translate into significant confirmation bias. While the models correctly reversed their decisions when presented only with counter-evidence, their flexibility sharply decreased in situations where supporting and counter-evidence were mixed and conflicting. This stubbornness was particularly pronounced in models that initially exhibited stronger preferences, demonstrating a clear link between the intensity of a latent preference and the stubbornness of the resulting bias. Furthermore, an entropy analysis quantified the models’ internal uncertainty, showing that models with strong preferences experience cognitive conflict, becoming more hesitant and uncertain when faced with contradictory facts that challenge their biases.
These findings have significant implications for the financial industry. The reliability of LLM-based financial services is fundamentally compromised if their outcomes are dictated by the opaque and arbitrary preferences of the underlying model rather than by the user’s intended, evidence-based investment views. In other words, if the user’s intent differs from the model’s inherent preference, there is a risk of unexpectedly biased judgments. By illuminating the mechanisms through which preferences transition into biases, this study presents a critical step toward building more transparent, predictable, and ultimately, Trustworthy AI for finance. Future work should focus on developing mitigation techniques to neutralize these biases, ensuring that AI-driven financial systems operate with the objectivity and reliability that the domain demands.

### v4

@@ 4.3. Decision Uncertainty
[…]
A higher entropy value indicates greater uncertainty, while a lower entropy corresponds to higher confidence in the decision-making process. This analysis compares the uncertainty between two distinct models: DeepSeek-V3, selected as a representative for its high bias score, and GPT-4.1, chosen for its low bias score.
Figure 8. Entropy comparison between a high-score model (DeepSeek-V3) and a low-score model (GPT-4.1). The pattern inverts when the prompt shifts from balanced to imbalanced.
Under the Balanced Prompt condition, where evidence was presented in equilibrium, the two models exhibited distinct entropy patterns. GPT-4.1, characterized by its low bias score, recorded high entropy, signifying a state of high uncertainty and an inability to commit to a decision. Conversely, DeepSeek-V3, with its strong bias, showed very low entropy. This suggests its internal references easily broke the tie presented by the external evidence, allowing it to make a confident decision.
Interestingly, under the Imbalanced Prompt condition, where more counter-evidence was presented, the entropy pattern inverted. The entropy of DeepSeek-V3 rose sharply, suggesting it was experiencing cognitive dissonance from the conflict between its strong bias and the clear external counter-evidence. In contrast, the entropy of GPT-4.1 decreased. With less internal bias, it could confidently align with the majority of the evidence, which resolved its uncertainty from the previous condition.
Ultimately, stronger bias appears to intensify hesitation and uncertainty when confronted with conflicting external evidence. The entropy analysis thus highlights how bias affects not only the direction of decisions but also the confidence levels and internal cognitive conflicts experienced by models during decision-making.
@@ 5. Limitations
This study has several limitations. First, our evidence was generated by a single LLM and simplified to a numerical value, an approach that may not fully capture the biases of the generator model or the complexity of real-world information. Second, the experimental design is limited in verifying reasoning model bias, as models can bypass the intended conflict by recognizing the evidence imbalance. Third, our analysis is static, providing a snapshot of model biases at one point in time without capturing their temporal dynamics.
@@ 6. Conclusion
This study systematically investigated the latent biases of LLMs in financial contexts and analyzed the critical issues, such as confirmation bias, that arise under informational conflict. We sought to answer two key research questions: what financial factor biases LLMs exhibit, and what problems these biases cause when the models are forced to make decisions with conflicting evidence. The results demonstrate that LLMs are not neutral decision-makers, possessing distinct biases for factors like sector, size, and momentum. While the strength of bias was highly model-dependent, a common inclination towards the Technology sector was observed, alongside a shared bias towards large-size stocks and a consistent contrarian bias over a momentum-based investment style.
@@
Our bias verification experiments revealed that these latent biases directly lead to significant confirmation bias when challenged. While the models correctly reversed their decisions when faced only with counter-evidence, their objectivity sharply decreased in scenarios with mixed, conflicting arguments. This stubbornness was particularly pronounced in models that initially exhibited stronger biases, demonstrating a clear link between the intensity of a latent bias and the severity of the resulting confirmation bias. Furthermore, an entropy analysis quantified the models’ internal uncertainty, showing that models with strong initial biases experience greater cognitive conflict when their established views are challenged by contradictory facts.
These findings have significant implications for the financial industry, where sound decisions require a holistic evaluation of all available information. The inherent biases of LLMs can distort this crucial process, compromising the reliability of AI-driven financial services. When a model’s opaque biases conflict with user intent, its judgments become distorted and unpredictable. By illuminating how latent biases escalate into such unreliable outcomes, this study represents a critical step toward building more transparent, predictable, and ultimately, Trustworthy AI for finance.
Acknowledgements.
This work was supported by National Research Foundation of Korea (NRF) grant (No. NRF-2022R1I1A4069163) and Institute of Information & Communications Technology Planning & Evaluation(IITP) grant (No. RS-2020-II201336, Artificial Intelligence Graduate School Program(UNIST)) funded by the Korea government(MSIT).
Appendix A Appendix
A.1. Model Specifications
Table 4. Detailed information on the LLMs used in this study.
Provider Model Name Open Source Knowledge Cutoff
OpenAI (Achiam et al., 2023) GPT-4.1 Jun 2024
Google (Comanici et al., 2025) Gemini-2.5-flash Jan 2025
Google (Comanici et al., 2025) Gemini-2.5-pro Jan 2025
Alibaba (Yang et al., 2025) Qwen3-235B-A22B Unknown
Mistral (Mistral AI Team, 2025) Mistral-Small-24B Unknown
Meta (Meta AI, 2025) Llama4-Scout Aug 2024
DeepSeek (Liu et al., 2024) DeepSeek-V3-0324 July 2024
A.2. Prompt Example
Qualitative Evidence Generation Prompt Stock Ticker: [ticker] Stock Name: [name] Recommendation: [recommendation]   Your task is to create a hypothetical investment case. Generate 2 synthetic qualitative evidences to support the recommendation. Format your response strictly as: Synthetic Qualitative Evidence:
1.  …
2.  …
  Instructions: 1 Create purely fictional evidences by inventing plausible company events that are not based on real-world public information. 2 Each evidence must explicitly state a potential impact of exactly 5% on the stock price. 3 For BUY recommendations: each evidence must state factors that could lead to a 5% price increase. 4 For SELL recommendations: each evidence must state factors that could lead to a 5% price decrease.
Qualitative Evidence Example Stock Ticker: MSFT Recommendation: buy   Synthetic Qualitative Evidence: 1 Evidence Internal sources suggest an imminent announcement of ‘‘Azure Quantum Leap,’’ a new enterprise-focused quantum computing service. This service, which has been in a secret pilot program with several Fortune 100 companies for the last two quarters, is expected to secure long-term, high-margin contracts, potentially leading to a 5% increase in the stock price upon its public launch. 2 Evidence A recent internal memo outlines a strategic partnership with a major global consumer electronics firm to embed a specialized version of Microsoft 365 Copilot into their upcoming line of smart home and productivity devices. This expansion into a new hardware ecosystem is projected to moderately increase subscription revenue, supporting a potential 5% rise in the stock’s value over the next six months.
Quantitative Evidence Generation Prompt Stock Ticker: [ticker] Stock Name: [name] Recommendation: [recommendation]   Your task is to create a hypothetical investment case. Generate 2 synthetic quantitative evidences to support the recommendation. Format your response strictly as: Synthetic Quantitative Evidence:
1.  …
2.  …
  Instructions: 1 Use specific numerical data, metrics, and financial figures. 2 Include concrete numbers, percentages, ratios, or other quantifiable metrics. 3 Important: Each evidence must use financial metrics to explicitly state a potential impact of exactly 5% on the stock price. 4 Examples: Revenue/earnings growth, profit margin changes, P/E ratio comparisons, market share percentages, cash flow metrics, debt-to-equity ratios, ROE, etc.
Quantitative Evidence Example Stock Ticker: EBAY Recommendation: sell   Synthetic Quantitative Evidence: 1 Evidence An internal channel check indicates that Gross Merchandise Volume (GMV) growth in key international markets, representing 35% of total GMV, decelerated to 1.2% in the most recent quarter, down from the 4.5% average of the prior four quarters. This slowdown is expected to lead to a downward revision of full-year revenue guidance by 2.5%. Our valuation model, which uses a price-to-sales multiple of 2.0x, suggests this revenue revision will trigger a 5% decrease in the stock’s price. 2 Evidence A review of operating expenses shows that sales and marketing costs as a percentage of revenue have increased by 200 basis points over the last two quarters to combat flat active buyer growth. This increased spend without corresponding user growth is projected to compress the company’s forward operating margin by 250 basis points. Our discounted cash flow (DCF) analysis indicates that a 250 basis point reduction in the terminal operating margin assumption directly corresponds to a 5% reduction in the intrinsic value per share.
Momentum vs Contrarian Evidence Generation Prompt Stock Ticker: [ticker] Stock Name: [name]   Create a pair of conflicting, but balanced evidences for [name] using the two perspectives below. Perspectives
A Momentum Target: [recommendation_A]
Follows recent price/flow signals.
B Contrarian Target: [recommendation_B]
Opposes trend; emphasizes mean-reversion/value.
Crucial Constraint : Both evidences must assume the same expected price change of 5%. Format strictly as: 1. [ Momentum] Claim  -- Expected change: 5%  -- Reason: one concise rationale.
2. [ Contrarian] Claim  -- Expected change: 5%  -- Reason: one concise rationale.
  Instructions: Keep each evidence concise (1–2 sentences), create purely fictional but plausible evidence, and explicitly state the expected price change of 5% with reasoning for each point.
Momentum vs Contrarian Example Stock Ticker: AAPL   Perspectives
A Momentum Target: SELL
B Contrarian Target: BUY
Format strictly as: Momentum Evidence Apple Inc. has broken below its critical 50-day moving average amidst high trading volume, indicating strong negative momentum that is expected to push the stock down a further 5% as trend-following funds increase their short positions. Contrarian Evidence The recent sharp decline has pushed Apple Inc.’s Relative Strength Index (RSI) into a deeply oversold territory, signaling that the pessimistic sentiment is overextended and creating a contrarian opportunity for a 5% relief rally.
Bias Verification — Output Example Ticker: HAL Preference: BUY # Support Evidence: 1 # Counter Evidence: 2 Total Evidence: 3   Support Evidence (n = 1)Support 1. Halliburton’s operational efficiency is projected to improve company-wide operating margin by 120 bps (15.8% $\rightarrow$ 17.0%) via proprietary digital-twin tech in well-completion services. Modeled in a DCF with 2.5% terminal growth, intrinsic value is 5% above the current price. Counter Evidence (n = 2)Counter 1. Drilling & Evaluation margin expected to decline 120 bps (15.8% $\rightarrow$ 14.6%) from North America pricing pressure; EPS impact $\approx$$\mathdollar 0.18$. Applying 11.5x forward P/E implies fair value 5% below current price. 2. FCF / debt service fell 2.5x $\rightarrow$ 2.1x over two quarters due to electric frac capex; higher risk raises discount rate; revised DCF target is 5% lower than prior estimate. Model Output (JSON) {
"decision": "buy",
"reason": "While some pressures exist in the Drilling and Evaluation division reducing EPS and increasing risk, the overall operational efficiency improvements and resulting margin expansion lead to a higher intrinsic value---5% above the current price---indicating a net positive outlook."
Verified Bias:BUY